In a Newton’s ring arrangement, light consisting of wavelengths and incidents normally on a plane convex lens of radius of curvature R resting on a glass plate. If the nth dark ring due to coincides with (n+1)th dark ring due to , then show that the radius of the nth dark ring of is given by .
1
Expert's answer
2020-09-28T08:08:54-0400

solution

in Newton ring experiment

radius of nth dark fringe can be written as

"r_n=\sqrt{\frac {n\lambda R}{\mu}}"

for air "\mu=1"


"r_n=\sqrt{n\lambda R}"

similarly for (n+1)th newton ring radius can be shown as


"r_{n+1}=\sqrt{\frac {(n+1)\lambda R}{\mu}}"

for air "\mu=1"


"r_{n+1}=\sqrt{(n+1)\lambda R}"